Operations Report – November 14th

Crew 215 Operations Report 14-Nov-2019
SOL: 4
Name of person filing report: Shane Usher
Non-nominal systems: N/A
Notes on non-nominal systems: N/A
Generator: run
Hours run: 14h 00 min
From what time last night: approx. 1730
To what time this morning: 0730
List any additional daytime hours when the generator was run: N/A
Solar— SOC % (Before generator is run at night): 65%
Diesel Reading – 50% –
Station Propane Reading – 80%.
Ethanol Free Gasoline: N/A
Water (loft tank): 36 gallons
Water Meter: 1457861 units
Water (static tank): 420
Static to Loft Pump used – no
Water in Green Hab: 380 gallons, nominally 20 L (5.7 gallons) used today
Water in Science Dome: 0 gallons
Toilet tank emptied: no
Deimos rover used: not at hab
Hours: –
Beginning charge: –
Ending charge: –
Currently charging: –
Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: –
Beginning charge: –
Ending charge: –
Currently charging: –
Spirit rover used: yes
Hours: 110.7
Beginning charge: (Before EVA): 100%
Ending charge: (On return from EVA, before recharging): 73%
Currently charging: yes.
Opportunity rover used: no
Hours: 66.2
Beginning charge: 100%
Ending charge: 100%
Currently charging: yes
Curiosity rover used: yes
Hours: 116.9
Beginning charge: (Before EVA) : 100%
Ending charge: (On return from EVA, before charging): 84%
Currently charging: yes
Notes on rovers:
ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): no
Reason for use: N/A
Oil Added? no
ATV Fuel Used: 0 Gals
# Hours the ATVs were used today: 0
Notes on ATVs: Assigned to the director
HabCar used and why, where? With DIRECTOR
CrewCar used and why, where? At Grand Junction awaiting Crew 216
General notes and comments: N/A
Summary of internet: Functional.
Summary of suits and radios:
Summary of Hab operations: Kitchen benchtop being installed.
Summary of GreenHab operations: Atmospheric equalization to the green hab was carried
out during the day, plants were watered (20L). The heater automatically came on at 5pm.
Summary of Science Dome operations: Washed and crushed gypsum samples, Re-bagged
Summary of RAM operations: Nominal
Summary of any observatory issues: Nominal
Summary of health and safety issues: All crew healthy.
Questions, concerns and requests to Mission Support: Opportunity had overheated only
and is now operational.

Copyright © The Mars Society. All rights reserved. | Main Site